OpenStudy (anonymous):

Determine the value of f(-6), f(3/2), f(2c) and f(c+1), then simplify as much as possible Question : f(x) = 3x^2 - 4x

OpenStudy (anonymous):

@Aylin help please :D

OpenStudy (anonymous):

Well, in the first two, just take the appropriate number and plug it in where x is. Then simplify it down. For the third one (2c) replace x with 2c, then see what you get. Same for the fourth one. The first two are plug'n'chug, but if you need help with the second two I'll stick around.

OpenStudy (anonymous):

the second two, just plug it in to every x in the problem ?

OpenStudy (anonymous):

Mhm. So f(2c) would like\[f(2c)=3 \times (2c)^{2}-4 \times (2c)\]

OpenStudy (anonymous):

Ok let me try thanks!

OpenStudy (anonymous):

ummm question

OpenStudy (anonymous):

Yes?

OpenStudy (anonymous):

kinda dumb question but for the first one f(-6) i did -6=3(-6)^2-4(-6) -6=108+24 -6=132 then at this step do you add the 6 over or minus the 132 over ?? lol

OpenStudy (anonymous):

it usually easier if one size has a (x)

OpenStudy (anonymous):

Well, the left side is f(-6) not -6. So when you get to f(-6)=132, you are finished with that part.

OpenStudy (anonymous):

Since you've found the value in the range (132) that corresponds to the value of -6 in the domain.

OpenStudy (anonymous):

i see

OpenStudy (anonymous):

the second one i found is 3/4 hope its right

OpenStudy (anonymous):

the third one i got: f(2c) = 4c

OpenStudy (anonymous):

and do you divide the 2c over ?

OpenStudy (anonymous):

I also got 3/4

OpenStudy (anonymous):

cool :D

OpenStudy (anonymous):

Well, with these problems you only need to simplify the right side, don't worry about the left side at all. In the third and fourth ones your answer will be in terms of c though.

OpenStudy (anonymous):

so is my answer for the third one okay ?

OpenStudy (anonymous):

Well, no.\[(2c)^{2}=4c^{2}\]and there is nothing in there to cancel it, so you should have a c^2 term.

OpenStudy (anonymous):

how did you get 2c^2 on the left side ?

OpenStudy (anonymous):

isn't it 2c ?

OpenStudy (anonymous):

\[f(2c)=3(2c)^{2}-4(2c)\]and\[(2c)^{2}=4c^{2}\]so\[f(2c)=3 \times 4c^{2}-4(2c)\]

OpenStudy (anonymous):

oh so you can't multiply the term -4 with 2c ?

OpenStudy (anonymous):

You can, but I was leaving it for you to do. :P

OpenStudy (anonymous):

ohhhh lol

OpenStudy (anonymous):

at first i thought u can make (2c)^2 = 4c and just X 3 = 12c

OpenStudy (anonymous):

i was wrong :|

OpenStudy (anonymous):

Pretty close, just forgot the ^2 part of it. :P

OpenStudy (anonymous):

u mean 12c^2 ?

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

but why ? :O isn't it already been done on the 2c^2 ?why ^2 another

OpenStudy (anonymous):

@___@

OpenStudy (anonymous):

Well\[(2c)^{2}\]is the same thing as\[2c \times 2c\]Does that make it more clear?

OpenStudy (anonymous):

does that make 4c ?

OpenStudy (anonymous):

and 4c x 3 = 12c lol ?

OpenStudy (anonymous):

\[2c \times 2c = 4c^{2}\]

OpenStudy (anonymous):

ohhh lol =___=

OpenStudy (anonymous):

omg...my mind is somewhere else today....sorry

OpenStudy (anonymous):

Mine is too, don't worry.

OpenStudy (anonymous):

naw, your good :D

OpenStudy (anonymous):

so once i got f(2c) = 12c^2 - 8c then what can i do next?

OpenStudy (anonymous):

Well, at this point all of the same powers of c are together, so you're done.

OpenStudy (anonymous):

what do you mean by "same powers of c are together" ? :d

OpenStudy (anonymous):

Well, there's only one c^2 term and only one c term. You can't combine the c^2 and the c terms, so there is nothing left to do.

OpenStudy (anonymous):

i see

OpenStudy (anonymous):

the fourth one im stuck completely :D

OpenStudy (anonymous):

Well, what is\[(c+1)^{2}\]?

OpenStudy (anonymous):

(c+1)(c+1)

OpenStudy (anonymous):

So FOIL it out! :P

OpenStudy (anonymous):

ohh let see

OpenStudy (anonymous):

c^2+2c+1 and i think u can factor it more right ?

OpenStudy (anonymous):

Well, we don't want to factor it right now. So now\[f(c+1)=3 \times (c^{2}+2c+1)-4 \times (c+1)\]Can you finish it from there?

OpenStudy (anonymous):

kk hold on

OpenStudy (anonymous):

i think it is : f(c+1)=3c^2 + 2c -1

OpenStudy (anonymous):

Yep!

OpenStudy (anonymous):

awesome! lol

OpenStudy (anonymous):

so is that it ? o_0

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

cool! thanks Aylin

OpenStudy (anonymous):

You're welcome.

OpenStudy (anonymous):

your really good in math :D genius maybe?

OpenStudy (anonymous):

Not really, I just want to teach math.

OpenStudy (anonymous):

btw are you in college?

OpenStudy (anonymous):

oh ! so your a teacher :D

OpenStudy (anonymous):

wow! how lucky i am :D

OpenStudy (anonymous):

Not yet. I'm a senior at RPI.

OpenStudy (anonymous):

But next year I hope to be a teacher.

OpenStudy (anonymous):

so your graduating next year ?

OpenStudy (anonymous):

In May

OpenStudy (anonymous):

oh wow, almost there!

OpenStudy (anonymous):

how many years did you spend in RPI ?

OpenStudy (anonymous):

The end of this year will make three years.

OpenStudy (anonymous):

that's really quick wow

OpenStudy (anonymous):

is this it ? "Rensselaer Polytechnic Institute" ?

OpenStudy (anonymous):

Yeah. Not really very fast though; I spent two years at another school first.

OpenStudy (anonymous):

I see, still not bad. I had a friend who spends 3 years at community college and transfer

OpenStudy (anonymous):

what's the GPA for acceptance in RPI?

OpenStudy (anonymous):

It was a community college that I came from, actually. From another college or from high school?

OpenStudy (anonymous):

another college

OpenStudy (anonymous):

yeah, everybody is doing community college first now aday, cuz it's so much cheaper :D

OpenStudy (anonymous):

Minimum is 3.0 Higher is better of course, but they also like it if you've been taking mostly science and math courses (or engineering if that's your major choice) if possible.

OpenStudy (anonymous):

make sense.

OpenStudy (anonymous):

well thanks so much for talking with me Aylin. I gotta get to dinner now :D

OpenStudy (anonymous):

Like if you took a lot of math/science courses and made mostly Bs, it would look better than if you had made As taking mostly things like Art or Chorus or History

OpenStudy (anonymous):

Bye

OpenStudy (anonymous):

anything with science and math is better...and it's the subjects that im most terrible at. WOW..

OpenStudy (anonymous):

well bye for now :d

OpenStudy (anonymous):

Well, the other thing is that RPI is about $60000 a year right now, so...

OpenStudy (anonymous):

your kidding me? no way

OpenStudy (anonymous):

http://admissions.rpi.edu/aid/index.html

OpenStudy (anonymous):

so how much for 1 semester of full time?

OpenStudy (anonymous):

omg wow...

OpenStudy (anonymous):

About half, but it depends on what your dining plan is, where you live, ect.

OpenStudy (anonymous):

omg so much debts...well but it's COLLEGE that we're talking about here so no suprise.. :|

OpenStudy (anonymous):

one time I read that article about this college couples graduated with over 200k debts scared the hell out of me :|

OpenStudy (anonymous):

I'm gonna have about $45K.

OpenStudy (anonymous):

did they tell you about the rate? or how much you have to pay back once your graduated ? I mean monthly?